OpenStudy (anonymous):

How do you factor x^2+8x-5?

OpenStudy (anonymous):

hint: use radicals and its conjugate...i.e., \[(\sqrt{a}+b)(\sqrt{a}-b) = a - b^2\]...let me know if you need more hints

OpenStudy (anonymous):

I'm not sure how to use that... I know the answer should be in (x +\- m)(x +\- n) format, but I can't seem to figure out how to get there.

OpenStudy (psymon):

Quadratic formula.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."